OpenStudy (bri-bri2017):

I have a homework paper about simplifing an expression and stating excluding values. I know how to simplify but how does it work with the excluding values? Thank you.

satellite73 (satellite73):

make sure the denominator is not zero

satellite73 (satellite73):

for example if you end up with \[\frac{1}{x-2}\] then you have to exclude \(x=2\)

satellite73 (satellite73):

also if you start with \[\frac{x+4}{(x-2)(x+4)}\] and cancel, then you have to exclude both \(2\) and \(-4\)

OpenStudy (bri-bri2017):

\[\frac{ 6p-36 }{ 18}\] That is one of my questions. I know it simplifies to p-6 over 3. How would I figure out the excluded value.

satellite73 (satellite73):

there are none since the denominator is a number

OpenStudy (bri-bri2017):

So if the equation is \[\frac{ 8b ^{5} }{ 64b^4 }\] it would be simplified to \[\frac{ 1 }{ 8 }b\] So they excluded would be b?

Vocaloid (vocaloid):

^ be careful. even though the simiplified version only has a number in the denominator, the original version excludes b = 0

OpenStudy (bri-bri2017):

Okay Thank you :)